I have a 30 word problem project due tomorrow. I've attempted doing the problem many times in different ways and neither of them give me the answer 100.
The problem is:
The length of a rectangular sheet of paper was twice its width. After 1 cm was trimmed from each edge of the sheet, the perimeter was at most 1 m. Find the largest possible dimensions of the trimmed sheet.
I've drawn this equation:
100 > 2(2w - 1) + 2(w - 1)
When I solved, I got w > 16
When i plug it into "l" I get 31.
my problem is that i 31 + 31 + 16 + 16 =/= 100
Please help ASAP!

Answer by mananth(16949)

width = w
length = 2w
You have missed the point that if you cut 1 cm from either side the length & width reduce by 2cm
w-2 new width
2w-2 new length
2(w-2)+2(2w-2)<100
2w-4+4w-4<100
6w-8<100
6w<108
/6
w<18 cm
w-2 = 16 maximum width
,,
2w-2= 34 maximum length
16+16+34+34= 100